undeadknight26 (undeadknight26):

Derive the equation of the parabola with a focus at (−5, 5) and a directrix of y = -1.

undeadknight26 (undeadknight26):

f(x) = -one twelfth (x - 5)2 + 2 f(x) = one twelfth (x - 5)2 + 2 f(x) = -one twelfth (x + 5)2 + 2 f(x) = one twelfth (x + 5)2 + 2